How to find extreme values?

1) P(z<z*)=0.05

2) P(z>z*)=0.025

3) find the "most extreme 5%" by separating it from the middle 95%.
Answer #1

1)

P(Z < z*) = 0.05

From Z table, z-score for the probability of 0.05 is -1.645

z* = -1.64

2 )

P(Z > z*) = 0.025

P(Z < z*) = 1 - 0.025

P(Z < z*) = 0.975

From Z table, z-score for the probability of 0.975 is 1.96

z* = 1.96

3)

We have to calculate z * such that P(-z* < Z < z*) = 0.95

P(-z* < Z < z*) = 0.95

P(Z < z*) - P(Z < -z*) = 0.95

P(Z < z*) - ( 1 - P(Z < z*) ) = 0.95

P(Z < z*) - 1 + P(Z < z*) = 0.95

2 P(Z < z*) = 1.95

P(Z < z*) = 0.975

From Z table, z-score for the probability of 0.975 is 1.96

z* = 1.96

( Two extreme values are -z* = -1.96 and z* = 1.96)